OpenStudy (anonymous):

How do you simplify: square root y^13

OpenStudy (anonymous):

y^(13/2)

OpenStudy (anonymous):

The answer was listed as y6 square root y. I am lost on how they got this.

OpenStudy (anonymous):

y^6square root y

OpenStudy (anonymous):

hm that doesnt make much sense the square and the square root cancel each other out. so in your answer you wouldnt have a radical (square root)

OpenStudy (anonymous):

We have this problem: \[\sqrt{y^{13}}\]Recall that square rooting something is equal to halving its exponent. In other words, we are raising the y^13 to the one half power. Therefore our answer will be: \[y^{13/2} = y^{6.5}\]Now, we can rewrite this using some basic exponents to get: \[y^{6.5}=y^6 \times y^{.5}=y^6 \times \sqrt{y}\]